OpenStudy (anonymous):

Algebra 1 help please! Evaluate 9 - 3 • 2 + 4. 6 7 15 16

OpenStudy (anonymous):

Recall your order of operations; multiplication has a higher precedence than addition or subtraction. Hence, this is the same as computing \(\large 9-(3\cdot 2)+4\); in other words, compute the product first and then combine everything else afterwards. Can you take things from here? :-)

OpenStudy (anonymous):

9 - 3 * 2 + 4 9 - 6 + 4 9 - 10 -1 That's what i keep getting

OpenStudy (anonymous):

Whenever you have a combination of + and -, you combine then going from left to right. This is important because 9 - 6 + 4 is not the same as 9 - (6 + 4), which is what you're doing (technically, 9 - (6 + 4) = 9 - 6 - 4, which is different from what you want to simplify originally). You want to treat 9 - 6 + 4 as (9 - 6) + 4 and then simplify that result instead. Does this clarify things? :-)

OpenStudy (anonymous):

So then the answer would be seven?

OpenStudy (anonymous):

That would be correct. :-)

OpenStudy (anonymous):

Can you see if i did this right?

OpenStudy (anonymous):

15 - 4 + (7 - 5)2 = __________. I got 7

OpenStudy (anonymous):

Just curious, is the last part supposed to read \((7-5)\cdot 2\) or \((7-5)^2\).?

OpenStudy (anonymous):

Oh sorry ^2

OpenStudy (anonymous):

Ok. We note that \((7-5)^2 = 2^2 = 4\). So it follows that \(15- 4 + (7-5)^2 = 15 - 4 + 4\). How should this simplify? Recall again that you want to start combining terms left to right. In other words, \(15 - 4 + 4 = (15-4)+4=\ldots\).

OpenStudy (anonymous):

So it would end up being fifthteen again?

OpenStudy (anonymous):

Yes, that's correct. Is this starting to make sense?

OpenStudy (anonymous):

Could you look at just one more please?

OpenStudy (anonymous):

Sure. :-)

OpenStudy (anonymous):

Evaluate 7 - (-1). 6 -6 8 -8

OpenStudy (anonymous):

I don't understand the extra symbols

OpenStudy (anonymous):

Do you know what the negative of a negative number is? (In other words, is the negative of a negative number positive or negative?) That's the key to solving this problem.

OpenStudy (anonymous):

Oh okay! So it's positive 6?

OpenStudy (anonymous):

No, that isn't correct. :-/ One thing that may help you with this is knowing that a negative number is the same as it's positive counterpart times -1; for example, \(-7 = (-1)\cdot 7\). So if you have \(-(-7)\), it's the same thing as saying \(-(-7) = (-1)\cdot(-7) = (-1)^2\cdot 7 = 7\). Hence, saying something like \(2-(-7)\) is equivalent to saying \(2+7\). Keeping this in mind, what would \(-(-1)\) be and thus, what is \(6-(-1)\)?

OpenStudy (anonymous):

Woops, that last statement should be \(7-(-1)\).

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

No. :-/ Let's take this step by step; in my example above, I showed you that \(-(-7)=7\). So what is \(-(-1)\)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Correct. So then it follows that \(7-(-1) = 7 + 1\). So what's the answer now? :-)

OpenStudy (anonymous):

8 :D

OpenStudy (anonymous):

Correct. Does this make more sense now? :-)

OpenStudy (anonymous):

Yeah, but if it isn't to much of a bother can we work on one more?

OpenStudy (anonymous):

Sure, that's fine with me. :-)

OpenStudy (anonymous):

Natalie made a mistake in the following problem. The mistake was made in Line ___. Line 1 6(4) ÷ 2 - 1 + 5 Line 2 6(4) ÷ 1 + 5 Line 3 24 ÷ 1 + 5 Line 4 24 + 5 Line 5 29

OpenStudy (anonymous):

I think she made a mistake in like 2

OpenStudy (anonymous):

That would be correct.

OpenStudy (anonymous):

What's the mistake? :-)

OpenStudy (anonymous):

cause when i tried to solve it i got 6

OpenStudy (anonymous):

she forgot to multiply

OpenStudy (anonymous):

That really isn't the mistake here. What changes from Line 1 to Line 2?

OpenStudy (anonymous):

the 2

OpenStudy (anonymous):

Yes. In particular she made the mistake thinking that \(6(4)\div 2 - 1 + 5 = 6(4)\div (2-1)+5\). Multiplication and division have higher precedence than addition and subtraction, so you need to simplify the multiplication and division first before doing any addition and subtraction.

OpenStudy (anonymous):

Thank you for all your help!! Have fun with you medal, new testimonial and a new fan :)!!!!

OpenStudy (anonymous):

It was a pleasure. I hope this order of operations business makes more sense now! :-)

OpenStudy (anonymous):

Yes it does!